The length of a rectangle if 4 feet less than 4 times the width. If the perimeter is 132, find the length and width of the rectangle.
Answer by rfer(16322)

2w+2l=P
2x+2(4x-4)=132
2x+8x-8=132
10x=140
x=14
4x-4=52